# On the Moreau envelope properties of weakly convex functions

### v1

@@ 1 Introduction
    $f(\lambda x+(1-\lambda)y)\leq\lambda f(x)+(1-\lambda)f(y),$ (1)
@@
Weakly convex functions, also called semi-convex functions [5, 19] play a central role in modern techniques for regularizing inverse problems with deep neural networks [9, 11, 17, 18, 22, 31, 33, 35, 36, 38]. Since they are non-convex, minimizing weakly convex functions remains a challenging task. A thorough understanding of their properties is therefore essential for effective optimization, especially when using proximal methods.
The Moreau envelope, also called Moreau-Yosida envelope, was first introduced in the seminal work of J.-J. Moreau [27] to regularize (non-differentiable) convex functions, thereby facilitating their optimization. In another context, such an envelope was introduced by K. Yosida in [40, Chapter IX, Section 4, page 240] to construct smooth approximations of the infinitesimal generator of a semi-group. Detailed explanation on the Moreau envelope for convex functions can be found in [34, Part 1.G].
The Moreau envelope regularizes in the sense that it is always differentiable (Proposition 5) and preserves the minimal value and minimizers of the original function (Proposition 6 and Lemma 13). However, it is not clear if the Moreau envelope "convexifies" the original function. Some results (Lemma 7 and Lemma 8) show that the Moreau envelope degrades both weak and strong convexity constants, whereas another one (Lemma 9) suggests that the Moreau envelope improves the convexity by non-increasing the non-convexity criteria (17).
All the results presented in this document are known by experts in the field. However, to our knowledge, there is no single reference that explicitly states and proves all of them, especially for weakly convex functions. Different proofs can be found in [1, 12, 16, 21, 27, 34]. Building on the work initiated in [33, Appendix D], we provide proofs that rely on elementary tools. For simplicity, we choose not to introduce the Clarke sub-differential [4] which can be used to analyze the Moreau envelope of non-differentiable functions. Nevertheless the results involving the differentiability assumption (Lemma 3(ii) and (iii), Proposition 4 and Lemma 14) can be generalized to non-differentiable functions using the Clarke sub-differential and following the same proof techniques. We refer to [20] for refined results on the regularity of the Moreau envelope in the non-convex and non-smooth setting and to [14] for a pedagogical introduction to the Moreau envelope in the convex case.
@@
First, in section 2, we define the Moreau envelope $g^{\gamma}$ and give some first immediate properties. In section 3, we analyze the dependency of $g^{\gamma}$ in the parameter $\gamma$. In section 4, we show the intrinsic link between the Moreau envelope and the convex conjugate, as well as a duality result (Proposition 3). In section 5, we investigate how the Moreau envelope preserves the convexity, the weak convexity and the strong convexity. In section 6, we derive the main properties of the proximal operator for weakly convex functions. In section 7, we prove that the Moreau envelope preserves critical points and minimizers, with the advantage of being differentiable. Finally, in section 8, the second derivative of the Moreau envelope is analyzed and we prove that the image of the proximal operator is almost convex.
@@ 2 Definitions and first properties
$f(\lambda x+(1-\lambda)y)\leq\lambda f(x)+(1-\lambda)f(y)+\frac{\rho}{2}\lambda(1-\lambda)\|x-y\|^{2}.$ (2)
@@
$\langle x-y,\nabla f(x)-\nabla f(y)\rangle\geq-\rho\|x-y\|^{2}.$ (3)
@@
$x^{T}\nabla^{2}f(y)x\geq-\rho\|x\|^{2}.$ (4)
Due to Definitions 1, note that convex functions are 0-weakly convex functions and $\mu$-strongly convex functions are $\mu-$weakly convex ones. Therefore, the classes of convex and strongly convex functions are included in the class of weakly convex functions, and the three above inequalities hold for $\rho=0$ (convex functions) or $\rho=-\mu$ ($\mu$-strongly convex functions).
@@
$g^{\gamma}(x)=\left(g\square\frac{1}{2\gamma}\|\cdot\|^{2}\right)(x)=\inf_{y\in{\mathbb{R}}^{d}}\frac{1}{2\gamma}\|x-y\|^{2}+g(y).$ (6)
@@
$\mathsf{Prox}_{\gamma g}\left(x\right)=\operatorname*{arg\,min}_{y\in{\mathbb{R}}^{d}}\frac{1}{2\gamma}\|x-y\|^{2}+g(y).$ (7)
@@
$g^{\gamma}(x)=g(\mathsf{Prox}_{\gamma g}\left(x\right))+\frac{1}{2\gamma}\|x-\mathsf{Prox}_{\gamma g}\left(x\right)\|^{2}.$ (8)
The Moreau envelope is an approximation of $g$ in the sense that $\forall x\in{\mathbb{R}}^{d}$, $\lim_{\gamma\to 0}g^{\gamma}(x)=g(x)$ (see Proposition 1) and it is a regularization as $g^{\gamma}$ is differentiable (see Proposition 5) even if $g$ is not.
@@
We will also show in Proposition 6 that $g$ and $g^{\gamma}$ have the same minimizers. Therefore, the Moreau envelope is particularly interesting for optimization. We can minimize the differentiable function $g^{\gamma}$ instead of the possibly non-differentiable $g$ [39].
@@
First, to analyze this dependency, we need a technical lemma, that will be useful for other properties (see Lemma 10).
Lemma 1.
@@
This lemma is a generalization of [1, Proposition 2.26] proposed in [16, Lemma 3.2].
Proof.
@@
$\lim_{\gamma\to 0}\mathsf{Prox}_{\gamma g}\left(x\right)=x,~~~~\lim_{\gamma\to 0}g(\mathsf{Prox}_{\gamma g}\left(x\right))=g(x).$ (10)
Proposition 1 shows that $g^{\gamma}$ has a monotonic dependency in $\gamma$. Moreover, when the step-size $\gamma\to 0$, the different objects have a well defined and computable limit.
Proof.
We need $\gamma<\frac{1}{\rho}$ to ensure that the Moreau envelope is well defined. We obtain from the definition of the Moreau envelope (see relation (6)) that $\gamma\to g^{\gamma}(x)$ is non-increasing and $g^{\gamma}(x)\leq g(x)$. The rest of the proof is a generalization of [1, Proposition 12.33] in the weakly convex case.
By applying Lemma 1, with $p=\textsf{Prox}_{\gamma g}(x)$, we have
@@
For $\frac{1}{\rho}>\mu\geq\gamma$, and $q=\mathsf{Prox}_{\mu g}\left(x\right)$. Lemma 1 gives that $\forall y\in{\mathbb{R}}^{d}$:
@@
Now, we prove that $g^{\gamma}(x)\to g(x)$ when $\gamma\to 0$. We denote $M=\sup_{\gamma\in(0,\frac{1}{2\rho}]}g^{\gamma}(x)<+\infty$. from the characterization (8) we have $\forall\gamma\in(0,\frac{1}{2\rho}]$,
@@
which proves the desired result. ∎
@@
Proposition 2 can be found in [13, Lemma 2 and Lemma 6]. Note that if $g$ is assumed to be continuous, we can prove that the function $(\gamma,x)\in[0,\frac{1}{\rho})\times{\mathbb{R}}^{d}\mapsto g^{\gamma}(x)$ is globally continuous by adapting the proof with the continuity extension theorem.
Proof.
@@
The previous inequality implies that $p=\mathsf{Prox}_{\gamma g}\left(x\right)$. So $\mathsf{Prox}_{\gamma g}\left(x\right)$ is the unique accumulation point of $\mathsf{Prox}_{\gamma_{n}g}\left(x_{n}\right)$, so we get that $\lim_{n\to+\infty}\mathsf{Prox}_{\gamma_{n}g}\left(x_{n}\right)=\mathsf{Prox}_{\gamma g}\left(x\right)$. This proves that $(\gamma,x)\in(0,\frac{1}{\rho})\times{\mathbb{R}}^{d}\mapsto\mathsf{Prox}_{\gamma g}\left(x\right)$ is continuous. Moreover, combining Proposition 1 and Lemma 11, we obtain that $\lim_{n\to+\infty}\mathsf{Prox}_{\gamma_{n}g}\left(x_{n}\right)=x$, if $x_{n}\to x$ and $\gamma_{n}\to 0$. So, we get that $(\gamma,x)\in[0,\frac{1}{\rho})\times{\mathbb{R}}^{d}\mapsto\mathsf{Prox}_{\gamma g}\left(x\right)$ is continuous.
@@ 4 The intrinsic link between the Moreau envelope and the convex conjugate
In this part, we introduce the convex conjugate and study its relation with the Moreau envelope. In particular, Proposition 3 shows a duality result between the convex conjugate and the Proximal operator.
Definition 4.
@@
The notion of convex conjugate of $f$ is related with the inf-convolution and the Moreau envelope, as shown by the following results from Lemma 2 to Proposition 3.
Lemma 2.
@@
Note that, by Lemma 2(i), $f^{\star}$ is convex even if $f$ is non-convex. Moreover, Lemma 2(ii) shows the compatibility of the convex conjugate with the inf-convolution operation.
Proof.
@@
Lemma 3(iii) can be reformulated as $\nabla f\circ\nabla f^{\star}=I_{d}$.
Proof.
@@
Thanks to Lemma 2(i) $f^{\star}$ is convex. Hence we get that $x=\nabla f^{\star}(y)$. The reverse is true because $f$ is convex so $f^{\star\star}=f$ by Lemma 2(iii). Thus, we can apply the same reasoning with $f^{\star}$ instead of $f$.
@@
$\langle\nabla f(x)-\nabla f(y),x-y\rangle\geq\mu\|x-y\|^{2}.$ (11)
@@
$f^{\star}(x)=\langle x,(\nabla f)^{-1}(x)\rangle-f((\nabla f)^{-1}(x)).$ (12)
@@
$\|(\nabla f)^{-1}(x)-(\nabla f)^{-1}(y)\|\leq\frac{1}{\mu}\|x-y\|.$ (13)
Now, we prove that $f^{\star}$ is differentiable and smooth. Using (12), we have for $x,h\in{\mathbb{R}}^{d}$
@@
Conversely, if $f^{\star}$ is $\frac{1}{\mu}$-smooth, since $f^{\star}$ is convex by the Baillon–Haddad theorem [1, Corollary 18.17], $\nabla f^{\star}$ is $\mu$-cocoercive, i.e. $\forall x,y\in{\mathbb{R}}^{d}$,
$\langle\nabla f^{\star}(x)-\nabla f^{\star}(y),x-y\rangle\geq\mu\|\nabla f^{\star}(x)-\nabla f^{\star}(y)\|^{2}.$ (14)
By denoting $z=\nabla f^{\star}(x)$ and $t=\nabla f^{\star}(y)$, by Lemma 3(iii), we have $x=\nabla f(z)$ and $y=\nabla f(t)$, so
@@
The previous result has been first proved in [27, Proposition 4.a]. Proposition 3 shows the intrinsic link between the proximal operator and the convex conjugate. In particular, it demonstrates that the inequality of Lemma 3(i) is attained only for pairs $(x,y)\in{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}$ that are of the form $(\mathsf{Prox}_{f}\left(z\right),\mathsf{Prox}_{f^{\star}}\left(z\right))$. Moreover, it also shows the equality $\mathsf{Prox}_{f}\left(z\right)+\mathsf{Prox}_{f^{\star}}\left(z\right)=z$.
Proof.
@@
which ends the proof. ∎

### v2

@@ Abstract
keywords:
Optimization, non convex, Moreau envelope
{MSCcodes}
49J52, 26B25, 46N10.
@@ 1 Introduction
    (1) $f(\lambda x+(1-\lambda)y)\leq\lambda f(x)+(1-\lambda)f(y),$
@@
Weakly convex functions have been first studied in [44], and are also called semi-convex functions [12, 33]. The class of weakly convex function is large as it includes convex functions and smooth functions, i.e. functions with Lipschitz gradient. More generally, the authors of [19, Lemma 4.2] show that any function of the form $h\circ c$, with $h$ convex and $c$ a smooth map with Lipschitz Jacobian is weakly convex. Note that such a composition is not necessary convex or smooth.
Many applied optimization problems can be reformulated using weakly convex functions, including phase retrieval [17], minimization of the conditional value-at-risk [54, 3], graph synchronization [1] or robust principal component analysis [6]. We refer to [18] for a survey on these applications. More recently, weakly convex functions play a central role in modern techniques for regularizing inverse problems with deep neural networks [21, 23, 31, 32, 36, 50, 52, 56, 57, 61]. Since they are non-convex, minimizing weakly convex functions remains a challenging task. A thorough understanding of their properties is therefore essential for effective optimization, especially when using proximal methods.
The Moreau envelope, also called Moreau-Yosida envelope, was first introduced in the seminal work of J.-J. Moreau [42] to regularize (non-differentiable) convex functions, thereby facilitating their optimization. In another context, such an envelope was introduced by K. Yosida in [63, Chapter IX, Section 4, page 240] to construct smooth approximations of the infinitesimal generator of a semi-group. Detailed explanation on the Moreau envelope for convex functions can be found in [55, Part 1.G]. The Moreau envelop has been used in many applications for its smoothing properties in optimization [4, 60, 29, 10] or in langevin diffusion [46, 20, 41, 52, 14, 25].
The Moreau envelope regularizes in the sense that it is always differentiable (Proposition 7.1) and preserves the minimal value and minimizers of the original function (Proposition 7.5 and Lemma 7.7). However, it is not clear if the Moreau envelope "convexifies" the original function. Some results (Lemma 5.7 and Lemma 5.9) indeed show that the Moreau envelope degrades both weak and strong convexity constants, whereas another one (Lemma 5.11) suggests that the Moreau envelope improves the convexity by non-increasing the non-convexity criteria (17).
The results presented in this document are known by experts in the field. However, to our knowledge, there is no single reference that explicitly states and proves all of them, especially for weakly convex functions. Different proofs can be found in [2, 24, 28, 35, 42, 55]. Building on the work initiated in [52, Appendix D], we provide here proofs that rely on elementary tools. For simplicity, we choose not to introduce the Clarke sub-differential [11] which can be used to analyze the Moreau envelope of non-differentiable functions. Nevertheless the results involving the differentiability assumption (Lemma 4.4(ii) and (iii), Proposition 6.1 and Lemma 7.9) can be generalized to non-differentiable functions using the Clarke sub-differential and following the same proof techniques. We also refer to [34] for refined results on the regularity of the Moreau envelope in the non-convex and non-smooth setting and to [26] for a pedagogical introduction to the Moreau envelope in the convex case.
@@
First, in section 2, we define the Moreau envelope $g^{\gamma}$ and give some first immediate properties. In section 3, we analyze the dependency of $g^{\gamma}$ in the parameter $\gamma$. In section 4, we show the intrinsic link between the Moreau envelope and the convex conjugate, as well as a duality result (Proposition 4.6). In section 5, we investigate how the Moreau envelope preserves the convexity, the weak convexity and the strong convexity. In section 6, we derive the main properties of the proximal operator for weakly convex functions. In section 7, we prove that the Moreau envelope preserves critical points and minimizers, with the advantage of being differentiable. Finally, in section 8, the second derivative of the Moreau envelope is analyzed and we prove that the image of the proximal operator is almost convex.
@@ 2 Definitions and first properties
(2) $f(\lambda x+(1-\lambda)y)\leq\lambda f(x)+(1-\lambda)f(y)+\frac{\rho}{2}\lambda(1-\lambda)\|x-y\|^{2}.$
@@
(3) $\langle x-y,\nabla f(x)-\nabla f(y)\rangle\geq-\rho\|x-y\|^{2}.$
@@
(4) $x^{T}\nabla^{2}f(y)x\geq-\rho\|x\|^{2}.$
Due to Definitions 1.1, note that convex functions are 0-weakly convex functions and $\mu$-strongly convex functions are $\mu-$weakly convex ones. Therefore, the classes of convex and strongly convex functions are included in the class of weakly convex functions, and the three above inequalities hold for $\rho=0$ (convex functions) or $\rho=-\mu$ ($\mu$-strongly convex functions).
@@
(6) $g^{\gamma}(x)=\left(g\square\frac{1}{2\gamma}\|\cdot\|^{2}\right)(x)=\inf_{y\in{\mathbb{R}}^{d}}\frac{1}{2\gamma}\|x-y\|^{2}+g(y).$
@@
(7) $\mathsf{Prox}_{\gamma g}\left(x\right)=\operatorname*{arg\,min}_{y\in{\mathbb{R}}^{d}}\frac{1}{2\gamma}\|x-y\|^{2}+g(y).$
@@
(8) $g^{\gamma}(x)=g(\mathsf{Prox}_{\gamma g}\left(x\right))+\frac{1}{2\gamma}\|x-\mathsf{Prox}_{\gamma g}\left(x\right)\|^{2}.$
The Moreau envelope is an approximation of $g$ in the sense that $\forall x\in{\mathbb{R}}^{d}$, $\lim_{\gamma\to 0}g^{\gamma}(x)=g(x)$ (see Proposition 3.3) and it is a regularization as $g^{\gamma}$ is differentiable (see Proposition 7.1) even if $g$ is not.
@@
We will also show in Proposition 7.5 that $g$ and $g^{\gamma}$ have the same minimizers. Therefore, the Moreau envelope is particularly interesting for optimization. We can minimize the differentiable function $g^{\gamma}$ instead of the possibly non-differentiable $g$ [62].
Example 2.3.
For $g(x)=-\frac{\rho}{2}\|x\|^{2}$ with $\rho\in{\mathbb{R}}$, and $\gamma>0$ such that $\gamma\rho<1$, we have $g^{\gamma}(x)=-\frac{\rho}{2(1-\gamma\rho)}\|x\|^{2}$.
In particular, if $\mu=-\rho>0$ and $g(x)=\frac{\mu}{2}\|x\|^{2}$, then $\forall\gamma>0$, we have $g^{\gamma}(x)=\frac{\mu}{2(1-\gamma\mu)}\|x\|^{2}$.
@@
First, to analyze this dependency, we need a technical lemma, that will be useful for other properties (see Lemma 6.3).
Lemma 3.1.
@@
This lemma is a generalization of [2, Proposition 2.26] proposed in [28, Lemma 3.2].
Proof 3.2.
@@
(10) $\lim_{\gamma\to 0}\mathsf{Prox}_{\gamma g}\left(x\right)=x,~~~~\lim_{\gamma\to 0}g(\mathsf{Prox}_{\gamma g}\left(x\right))=g(x).$
Proposition 3.3 shows that $g^{\gamma}$ has a monotonic dependency in $\gamma$. Moreover, when the step-size $\gamma\to 0$, the different objects have a well defined and computable limit. Note that in particular, we have $\forall\gamma\in(0,\frac{1}{\rho}),\forall x\in{\mathbb{R}}^{d},\inf g\leq g^{\gamma}(x)\leq g(x)$.
Proof 3.4.
We need $\gamma<\frac{1}{\rho}$ to ensure that the Moreau envelope is well defined. We obtain from the definition of the Moreau envelope (see relation (6)) that $\gamma\to g^{\gamma}(x)$ is non-increasing and $g^{\gamma}(x)\leq g(x)$. The rest of the proof is a generalization of [2, Proposition 12.33] in the weakly convex case.
By applying Lemma 3.1, with $p=\textsf{Prox}_{\gamma g}(x)$, we have
@@
For $\frac{1}{\rho}>\mu\geq\gamma$, and $q=\mathsf{Prox}_{\mu g}\left(x\right)$. Lemma 3.1 gives that $\forall y\in{\mathbb{R}}^{d}$:
@@
We now show that $g^{\gamma}(x)\to g(x)$ when $\gamma\to 0$. We denote $M=\sup_{\gamma\in(0,\frac{1}{2\rho}]}g^{\gamma}(x)<+\infty$. From the characterization (8) we have $\forall\gamma\in(0,\frac{1}{2\rho}]$,
@@
which proves the limit when $\gamma\to 0$.
@@
Proposition 3.5 can be found in [25, Lemma 2 and Lemma 6]. Note that if $g$ is assumed to be continuous, we can prove that the function $(\gamma,x)\in[0,\frac{1}{\rho})\times{\mathbb{R}}^{d}\mapsto g^{\gamma}(x)$ is globally continuous by adapting the proof with the continuity extension theorem.
Proof 3.6.
@@
The previous inequality implies that $p=\mathsf{Prox}_{\gamma g}\left(x\right)$. So $\mathsf{Prox}_{\gamma g}\left(x\right)$ is the unique accumulation point of $\mathsf{Prox}_{\gamma_{n}g}\left(x_{n}\right)$, so we get that $\lim_{n\to+\infty}\mathsf{Prox}_{\gamma_{n}g}\left(x_{n}\right)=\mathsf{Prox}_{\gamma g}\left(x\right)$. This proves that $(\gamma,x)\in(0,\frac{1}{\rho})\times{\mathbb{R}}^{d}\mapsto\mathsf{Prox}_{\gamma g}\left(x\right)$ is continuous. Moreover, combining Proposition 3.3 and Lemma 6.5, we obtain that $\lim_{n\to+\infty}\mathsf{Prox}_{\gamma_{n}g}\left(x_{n}\right)=x$, if $x_{n}\to x$ and $\gamma_{n}\to 0$. So, we get that $(\gamma,x)\in[0,\frac{1}{\rho})\times{\mathbb{R}}^{d}\mapsto\mathsf{Prox}_{\gamma g}\left(x\right)$ is continuous.
@@ 4 The intrinsic link between the Moreau envelope and the convex conjugate
In this part, we introduce the convex conjugate and study its relation with the Moreau envelope. In particular, Proposition 4.6 shows a duality result between the convex conjugate and the Proximal operator.
Definition 4.1.
@@
The notion of convex conjugate of $f$ is related with the inf-convolution and the Moreau envelope, as shown by the following results from Lemma 4.2 to Proposition 4.6.
Lemma 4.2.
@@
By Lemma 4.2(i), $f^{\star}$ is convex even if $f$ is non-convex. Moreover, Lemma 4.2(ii) shows the compatibility of the convex conjugate with the inf-convolution operation.
Proof 4.3.
@@
Lemma 4.4(iii) can be reformulated as $\nabla f\circ\nabla f^{\star}=I_{d}$.
Proof 4.5.
@@
Thanks to Lemma 4.2(i) $f^{\star}$ is convex. Hence we get that $x=\nabla f^{\star}(y)$. The reverse is true because $f$ is convex so $f^{\star\star}=f$ by Lemma 4.2(iii). Thus, we can apply the same reasoning with $f^{\star}$ instead of $f$.
@@
(11) $\langle\nabla f(x)-\nabla f(y),x-y\rangle\geq\mu\|x-y\|^{2}.$
@@
(12) $f^{\star}(x)=\langle x,(\nabla f)^{-1}(x)\rangle-f((\nabla f)^{-1}(x)).$
@@
(13) $\|(\nabla f)^{-1}(x)-(\nabla f)^{-1}(y)\|\leq\frac{1}{\mu}\|x-y\|.$
Now, we prove that $f^{\star}$ is differentiable and smooth.
Using (12), we have for $x,h\in{\mathbb{R}}^{d}$
@@
Conversely, if $f^{\star}$ is $\frac{1}{\mu}$-smooth, since $f^{\star}$ is convex by the Baillon–Haddad theorem [2, Corollary 18.17], $\nabla f^{\star}$ is $\mu$-cocoercive, i.e. $\forall x,y\in{\mathbb{R}}^{d}$,
(14) $\langle\nabla f^{\star}(x)-\nabla f^{\star}(y),x-y\rangle\geq\mu\|\nabla f^{\star}(x)-\nabla f^{\star}(y)\|^{2}.$
By denoting $z=\nabla f^{\star}(x)$ and $t=\nabla f^{\star}(y)$, by Lemma 4.4(iii), we have $x=\nabla f(z)$ and $y=\nabla f(t)$, so
@@
The previous result has been first proved in [42, Proposition 4.a]. Proposition 4.6 shows the intrinsic link between the proximal operator and the convex conjugate. In particular, it demonstrates that the inequality of Lemma 4.4(i) is attained only for pairs $(x,y)\in{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}$ that are of the form $(\mathsf{Prox}_{f}\left(z\right),\mathsf{Prox}_{f^{\star}}\left(z\right))$. Moreover, it also shows the equality $\mathsf{Prox}_{f}\left(z\right)+\mathsf{Prox}_{f^{\star}}\left(z\right)=z$.
Proof 4.7.
@@
which ends the proof.